# A convex formulation of covariate-adjusted Gaussian graphical models via natural parametrization

### v1

@@ Abstract
Gaussian graphical models (GGMs) are widely used for recovering the conditional independence structure among random variables. Recently, several key advances have been made to exploit an additional set of variables for better estimating the GGMs of the variables of interest. For example, in co-expression quantitative trait locus (eQTL) studies, both the mean expression level of genes as well as their pairwise conditional independence structure may be adjusted by genetic variants local to those genes. Existing methods to estimate covariate-adjusted GGMs either allow only the mean to depend on covariates or suffer from poor scaling assumptions due to the inherent non-convexity of simultaneously estimating the mean and precision matrix. In this paper, we propose a convex formulation that jointly estimates the covariate-adjusted mean and precision matrix by utilizing the natural parametrization of the multivariate Gaussian likelihood. This convexity yields theoretically better performance as the sparsity and dimension of the covariates grow large relative to the number of samples. We verify our theoretical results with numerical simulations and perform a reanalysis of an eQTL study of glioblastoma multiforme (GBM), an aggressive form of brain cancer.
@@ 1 Introduction
Graphical models are used to represent the distribution of a random vector $\bm{X}=(X_{1},\dotsc,X_{p})$ by relating its conditional independence structure to a graph. This correspondence is particularly salient when $\bm{X}$ is Gaussian. Letting $\boldsymbol{\mathbf{\Omega}}=\boldsymbol{\mathbf{\Sigma}}^{-1}$ be the precision matrix of a Gaussian random vector $\bm{X}$, for $i\neq j$, the components $X_{i}$ and $X_{j}$ are conditionally independent given $\bm{X}_{\{1,\dotsc,p\}\setminus{\{i,j\}}}$ if and only if $\Omega_{ij}=0$ [Lauritzen, 1996]. Models that estimate the conditional independence structure by imposing sparsity on $\boldsymbol{\mathbf{\Omega}}$ are known as Gaussian graphical models (GGMs).
There has been considerable research on GGMs and their estimation in the high-dimensional setting [Yuan and Lin, 2007, Friedman et al., 2007, Meinshausen and Bühlmann, 2006]. GGMs find major applications in genomics where the network structure of genes is of interest. Such a setting is typically high-dimensional such that the number of genes exceeds the number of samples [Schäfer and Strimmer, 2005]. Of particular interest in genomics are co-expression quantitative trait locus (eQTL) studies. In these studies, gene expression data are analyzed alongside information about external genetic markers that are known to confound the expression levels of genes. In order to use GGMs for eQTL studies, one must allow the GGM framework to incorporate covariate information.
Let $\bm{X}\in\mathbb{R}^{p}$ be the vector of responses and $\bm{U}\in\mathbb{R}^{q}$ the associated vector of covariates. We consider a general model
$\bm{X}\>|\>\bm{U}=\boldsymbol{\mathbf{u}}\sim N(\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}}),\boldsymbol{\mathbf{\Omega}}^{-1}(\boldsymbol{\mathbf{u}}))$ (1)
and a specification given by
@@
Note that in this specification, the error covariance matrix does not depend on covariates. Multivariate regression models specify (2) with the goal to improve estimation of $\boldsymbol{\mathbf{\Gamma}}$ by also estimating $\boldsymbol{\mathbf{B}}_{0}$; see for example Yuan et al. [2007] and references therein for this well-studied setting. By contrast, the focus of the covariate-adjusted GGM framework is to estimate the sparsity pattern of the precision matrix while accounting for covariate information through $\boldsymbol{\mathbf{B}}_{0}$. Several methods exist in this framework that employ sparse penalization to the negative log-likelihood of (1) [Rothman et al., 2010, Yin and Li, 2011, Cai et al., 2013, Chen et al., 2016, Wang, 2015, Chen et al., 2018]. These models also are termed “conditional Gaussian graphical models” in Yin and Li [2011].
Estimating (1) with (2) is challenging because the negative log-likelihood of (1) is not jointly convex in $\boldsymbol{\mathbf{\Gamma}}$ and $\boldsymbol{\mathbf{B}}_{0}$. Therefore, covariate-adjusted GGM methods use either alternating [Rothman et al., 2010, Yin and Li, 2011, Chen et al., 2018] or two-stage [Cai et al., 2013, Chen et al., 2016] estimation algorithms. Despite these challenges, minimax-optimal error rates in jointly estimating (2) have been established [Chen et al., 2018, Lv et al., 2022].
Although there has been progress in allowing heterogeneous means in GGMs, there is little work on allowing the graph structure of $\boldsymbol{\mathbf{\Omega}}$ to also depend on covariates; indeed the aforementioned methods all assume that $\boldsymbol{\mathbf{B}}_{0}$ is fixed across subjects. Yet subject-specific graph structures are plausible in applications such as eQTL analyses. For example, gene C may mediate the co-expression of genes A and B, but only in the presence of a single-nucleotide polymorphism (SNP) local to gene C. In other words, the expression of genes A and B may be conditionally independent given the rest of the network unless a genetic variant is present near gene C [Fehrmann et al., 2011, Kolberg et al., 2020, Rockman and Kruglyak, 2006]. To capture this additional structure, Zhang and Li [2022] extends (2) to allow both the mean and the network structure of the responses to be adjusted by covariates. Their model specifies (1) by taking
@@
where $\boldsymbol{\mathbf{B}}_{h},h=0,1\dotsc,q$ are sparse, symmetric matrices. Like previous works in covariate-adjusted GGMs, the joint estimation of the parameters $(\boldsymbol{\mathbf{\Gamma}},\boldsymbol{\mathbf{B}}_{0},\boldsymbol{\mathbf{B}}_{1},\dotsc,\boldsymbol{\mathbf{B}}_{h})$ in (3) involves a non-jointly-convex objective. Therefore Zhang and Li [2022] use a two-stage estimation procedure like in preceding methods.
@@
In this work, we contribute a jointly convex formulation of (1) such that both the mean and the graph structure may depend on covariates. Like Zhang and Li [2022], our method uses nodewise regression to estimate the covariate-adjusted graph structure [Meinshausen and Bühlmann, 2006]. However, we base our formulation upon a natural parametrization of the multivariate Gaussian likelihood such that each nodewise regression is a convex optimization problem.
1.1 Previous work
Our method is inspired by the penalized likelihood-based method of Zhang and Li [2022]. Bayesian approaches to subject-level precision matrix adjustment have been considered before. In Niu et al. [2024], a Bayesian method is developed where a discrete set of precision matrices $\boldsymbol{\mathbf{\Omega}}_{1},\dotsc,\boldsymbol{\mathbf{\Omega}}_{K}$ is determined for a partition of $n$ samples into $K$ sets, with the partition depending on external covariates. However, this approach does not immediately allow for a covariate-adjusted mean, nor is it easy to interpret how each covariate affects the overall graph structure through the $\boldsymbol{\mathbf{\Omega}}_{i}$. Wang et al. [2022] develops a Bayesian approach to solve similar nodewise regression problems as Zhang and Li [2022], utilizing sparsity-inducing priors to estimate the sparse component matrices. Their method also does not allow a covariate-adjusted mean, nor does it identify group structures.
Convex formulations of covariate-adjusted GGMs have been proposed before. Wang [2015] remarks on the non-convex nature of existing covariate-adjusted GGM methods and suggests a convex formulation of (1) with (2). However, their method achieves convexity by relying on an initial estimate of $\boldsymbol{\mathbf{\Gamma}}$, thereby placing it in the category of two-stage methods. A related line of work is the conditional Gaussian random field, which is also sometimes called the conditional Gaussian graphical model [Sohn and Kim, 2012, Yuan and Zhang, 2014]. This class of methods investigates estimating a sparse $\boldsymbol{\mathbf{B}}_{0}$ in (1) and (2) by penalizing a convex negative log-likelihood. This parametrization is made possible by the additional assumption that $(\bm{X},\bm{U})$ is jointly Gaussian. In Zhu [2020], the matrix $\boldsymbol{\mathbf{\Gamma}}$ in (2) is written as $\boldsymbol{\mathbf{\Gamma}}=\tilde{\boldsymbol{\mathbf{\Gamma}}}\boldsymbol{\mathbf{\Omega}}$ leading to a negative log-likelihood of $\boldsymbol{\mathbf{X}}$ that is jointly convex in $\tilde{\boldsymbol{\mathbf{\Gamma}}}$ and $\boldsymbol{\mathbf{\Omega}}$. However, their method does not immediately allow for covariate-adjusted network structures.
Our proposed method differs from these works as it allows both the mean and the precision matrix to depend on external covariates while also offering the advantage of solving a convex optimization problem.
1.2 Outline
We recall the Gaussian graphical regression framework of Zhang and Li [2022] and motivate the natural parametrization in Section 2. In Section 3 we provide the model specification and algorithm to induce a sparse-group structure in the graph. Our theoretical results are discussed in Section 4, namely that the natural parametrization allows for better theoretical scaling of $p$ and $q$ relative to $n$ under the same assumptions as in Zhang and Li [2022]. This is demonstrated through extensive simulations in Section 5. Finally we apply our method to perform a reanalysis of data from glioblastoma microforme (GBM) tissue samples in Section 6.
2 Parametrizations for Covariate-adjusted Graphical Models
Let $\bm{X}\in\mathbb{R}^{p}$ be a random vector of responses and $\bm{U}\in\mathbb{R}^{q}$ the corresponding covariates. In Zhang and Li [2022], the dependence of $\bm{X}$ on $\bm{U}$ is given by (1) and (3) so that both the mean and precision matrix are covariate-dependent. Write $\boldsymbol{\mathbf{\Gamma}}=(\boldsymbol{\mathbf{\gamma}}_{1},\dotsc,\boldsymbol{\mathbf{\gamma}}_{p})^{\top}$ and let $\Omega_{jk}(\boldsymbol{\mathbf{u}})$ denote the $(j,k)$-th element of $\boldsymbol{\mathbf{\Omega}}(\boldsymbol{\mathbf{u}})$, keeping in mind its dependence on $\boldsymbol{\mathbf{u}}$. Then (3) may be estimated via a neighborhood regression method [Meinshausen and Bühlmann, 2006], which amounts to estimating $p$ separate linear regression models of the form
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\sum_{k\neq j}^{p}\beta_{jk0}(X_{k}-\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{k})+\sum_{k\neq j}^{p}\sum_{h=1}^{q}\beta_{jkh}u_{h}(X_{k}-\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{k})+\varepsilon_{j}$ (4)
where $\beta_{jkh}=-[\boldsymbol{\mathbf{B}}_{h}]_{jk}/\Omega_{jj}$ and $\varepsilon_{j}\sim N(0,1/\Omega_{jj})$ for all $j$, $k$, and $h$. This is termed Gaussian graphical regression in Zhang and Li [2022].
Note that a least squares criterion based on (4) is not jointly convex owing to the cross term $\beta_{jkh}\times\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{k}$. Hence, Zhang and Li [2022] uses a two-stage estimation method. First, $\boldsymbol{\mathbf{\gamma}}_{j}$ is estimated for all $j$ via the model
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\xi_{j},\quad\operatorname{\operatorname{\mathbb{E}}}\xi_{j}=0,$ (5)
and $\ell_{1}$ penalization is added so that $\hat{\boldsymbol{\mathbf{\gamma}}}_{j}$ is sparse. Second, the observed response vectors $\boldsymbol{\mathbf{x}}_{j}$ are centered using $\hat{\boldsymbol{\mathbf{\gamma}}}_{j}$ and the coefficients $\beta_{jkh}$ in (4) are estimated with $\hat{\boldsymbol{\mathbf{\gamma}}}_{j}$ in place of $\boldsymbol{\mathbf{\gamma}}_{j}$ for all $j\in[p]$. As with other two-stage methods, the above procedure incurs an estimation error in the first stage because (5) ignores the dependence on the $\beta_{jkh}$ terms in (4). Assumptions on the scaling of the ambient dimensions $p$ and $q$ and the sparsity of the coefficient vectors are needed to suppress the model misspecification errors. Next, we will present a formulation of (1) so that the corresponding least squares criterion is convex, obviating the need for a two-stage procedure.
2.1 Convex formulation
Recall the form of the $p$-dimensional multivariate Gaussian likelihood expressed in terms of the natural parameters $(\tilde{\boldsymbol{\mathbf{\theta}}},\tilde{\boldsymbol{\mathbf{\Theta}}})$ where $\tilde{\boldsymbol{\mathbf{\theta}}}=\boldsymbol{\mathbf{\Omega}}\boldsymbol{\mathbf{\mu}}$ and $\tilde{\boldsymbol{\mathbf{\Theta}}}=-\boldsymbol{\mathbf{\Omega}}$:
$L(\tilde{\boldsymbol{\mathbf{\theta}}},\tilde{\boldsymbol{\mathbf{\Theta}}}\>|\>\boldsymbol{\mathbf{x}})=\exp\left\{\tilde{\boldsymbol{\mathbf{\theta}}}^{\top}\boldsymbol{\mathbf{x}}-\frac{1}{2}\boldsymbol{\mathbf{x}}^{\top}\tilde{\boldsymbol{\mathbf{\Theta}}}\boldsymbol{\mathbf{x}}-A(\tilde{\boldsymbol{\mathbf{\theta}}},\tilde{\boldsymbol{\mathbf{\Theta}}})\right\}.$
Our formulation is motivated by the fact that the cumulant function $A$ is jointly convex in $(\tilde{\boldsymbol{\mathbf{\theta}}},\tilde{\boldsymbol{\mathbf{\Theta}}})$ and therefore so is $-\log L(\tilde{\boldsymbol{\mathbf{\theta}}},\tilde{\boldsymbol{\mathbf{\Theta}}}\>|\>\boldsymbol{\mathbf{x}})$. Define

### v2

# Convex estimation of Gaussian graphical regression models with covariates

@@ Abstract
Gaussian graphical models (GGMs) are widely used to recover the conditional independence structure among random variables. Recent work has sought to incorporate auxiliary covariates to improve estimation, particularly in applications such as co-expression quantitative trait locus (eQTL) studies, where both gene expression levels and their conditional dependence structure may be influenced by genetic variants. Existing approaches to covariate-adjusted GGMs either restrict covariate effects to the mean structure or lead to nonconvex formulations when jointly estimating the mean and precision matrix. In this paper, we propose a convex framework that simultaneously estimates the covariate-adjusted mean and precision matrix via a natural parametrization of the multivariate Gaussian likelihood. The resulting formulation enables joint convex optimization and yields improved theoretical guarantees under high-dimensional scaling, where the sparsity and dimension of covariates grow with the sample size. We support our theoretical findings with numerical simulations and demonstrate the practical utility of the proposed method through a reanalysis of an eQTL study of glioblastoma multiforme (GBM), an aggressive form of brain cancer.
@@ 1 Introduction
Graphical models use graphs to represent conditional independence relationships among components of a random vector $\bm{X}=(X_{1},\dots,X_{p})$. In the Gaussian case, this leads to Gaussian graphical models (GGMs), where conditional independences correspond to zeros in the precision matrix $\boldsymbol{\mathbf{\Omega}}=\boldsymbol{\mathbf{\Sigma}}^{-1}$. Specifically, for a Gaussian random vector $\bm{X}$ and $i\neq j$, the variables $X_{i}$ and $X_{j}$ are conditionally independent given $\bm{X}_{\{1,\dots,p\}\setminus\{i,j\}}$ if and only if $\Omega_{ij}=0$ [lauritzen]. There has been extensive work on the estimation of GGMs in high-dimensional settings [yuanlin, friedman-glasso, meinshausen], where sparsity is typically induced on the precision matrix to improve interpretability and estimation accuracy.
A major application of GGMs arises in genomics, where the goal is to infer gene interaction networks from gene expression data. Such problems are inherently high-dimensional, with the number of genes often exceeding the number of samples [schafer2005]. A prominent example is provided by co-expression quantitative trait locus (eQTL) studies, in which gene expression measurements are analyzed jointly with external genetic markers. These markers may confound gene expression levels, either through local genetic effects or through associations with unmeasured factors influencing gene expression. This setting motivates the development of Gaussian graphical models that incorporate covariate information, enabling adjustment for such confounding in network estimation.
Specifically, let $\bm{X}\in\mathbb{R}^{p}$ be the vector of responses and $\bm{U}\in\mathbb{R}^{q}$ the associated vector of covariates. We consider the general model
$\bm{X}\>|\>\bm{U}=\boldsymbol{\mathbf{u}}\sim N(\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}}),\boldsymbol{\mathbf{\Omega}}^{-1}(\boldsymbol{\mathbf{u}})),$ (1)
which allows both the mean vector and the precision matrix of $\bm{X}$ to depend on a realization of covariates $\bm{U}=\boldsymbol{\mathbf{u}}$. A commonly used specification assumes that only the mean depends on $\bm{U}$, while the precision matrix remains constant. This specification is also known as the conditional Gaussian graphical model [yin-li]:
@@
Estimation of the unknown parameters in (2) is challenging, as the negative log-likelihood associated with (1) is not jointly convex in $(\boldsymbol{\mathbf{\Gamma}},\boldsymbol{\mathbf{B}}_{0})$. Existing approaches typically rely on alternating optimization [rothman, yin-li, chen18] or two-stage procedures [cai13, chen2016]. Despite this nonconvexity, minimax-optimal rates for joint estimation under (1) with (2) have been established [chen18, lv]. Several works have sought to convexify the problem. For instance, wang_joint_2015 propose a convex formulation based on an initial estimate of $\boldsymbol{\mathbf{\Gamma}}$, while zhu_convex_2020 introduce a reparameterization that yields a jointly convex negative log-likelihood.
Model (1) with specification (2) can also be viewed as a multivariate response regression model, where the primary goal is to improve estimation of $\boldsymbol{\mathbf{\Gamma}}$ by leveraging structural assumptions on $\boldsymbol{\mathbf{B}}_{0}$. In contrast, the covariate-adjusted GGM framework focuses on recovering the sparsity pattern of the precision matrix while accounting for covariate effects through $\bm{U}$. Several methods in this setting incorporate structure-inducing penalties into the likelihood of (1) associated with the specification (2); see liu_estimation_2025 for a recent review highlighting connections and distinctions between these closely related lines of work.
While substantial progress has been made in modeling heterogeneous means, relatively little work allows the graph structure itself to vary with covariates; most existing approaches instead assume a common precision matrix $\boldsymbol{\mathbf{B}}_{0}$ across subjects. However, subject-specific graph structures are plausible in applications such as eQTL studies. For example, gene C may mediate the co-expression of genes A and B only in the presence of a single-nucleotide polymorphism (SNP) local to gene C. In such cases, genes A and B may be conditionally independent given the rest of the network unless a genetic variant is present near gene C [fehrmann, kolberg, rockman].
A small number of works have considered models in which both the mean and precision depend on covariates. Bayesian approaches, for example, allow subject-specific precision matrices by imposing regression or clustering structures on $\boldsymbol{\mathbf{\Omega}}$ [ni_bayesian_2022, niu_covariate-assisted_2024, wang_bayesian_2022, zeng_bayesian_2025]. However, these methods do not directly address joint estimation of both the mean and precision as functions of covariates within a unified framework.
This work is closely related to zhangli, which extends (2) to allow both the mean and the network structure of the responses to depend on covariates. Specifically, their model specifies (1) by taking
@@
where $\boldsymbol{\mathbf{B}}_{h}\in\mathbb{R}^{p\times p},h=0,1\dotsc,q$ are sparse, symmetric matrices. As in conditional GGMs, joint estimation of the parameters $(\boldsymbol{\mathbf{\Gamma}},\boldsymbol{\mathbf{B}}_{0},\boldsymbol{\mathbf{B}}_{1},\dotsc,\boldsymbol{\mathbf{B}}_{q})$ under (3) leads to a nonconvex objective. To address this, zhangli adopt a two-stage estimation procedure. Building on this framework, zhang-multi-task-2025 propose a multi-task estimator with a cross-task group penalty, while meng_statistical_2025 develop debiased nodewise estimators for inference on edge–covariate effects.
@@
In this work, we develop Covariate-adjusted Sparse Precision with Natural Estimation (cspine), a method for estimating covariate-adjusted GGMs with heterogeneous network effects based on a jointly convex formulation of (1). Like zhangli, our method uses nodewise regression [meinshausen] to estimate the covariate-adjusted graph structure. However, by adopting a natural parametrization of the Gaussian likelihood, we obtain a formulation in which each nodewise regression reduces to a convex optimization problem, thereby avoiding the nonconvexity inherent in the existing approaches.
Our main contributions and the organization of the paper are as follows. In Section 2, we review the framework of zhangli and introduce the proposed natural parametrization, under which the likelihood in (1) is jointly convex in all unknown parameters. Building on this formulation, Section 3 presents the proposed model and estimation procedure with sparsity-inducing penalties. Section 4 establishes theoretical guarantees for the proposed estimator, showing that the convex formulation leads to improved scaling of the ambient dimensions $p$ and $q$ relative to the sample size $n$ under assumptions comparable to those in zhangli. These theoretical advantages are further supported by extensive simulations in Section 5. Finally, in Section 6, we illustrate the practical utility of the proposed approach through an analysis of glioblastoma multiforme (GBM) gene expression data. Proofs of all theoretical results, along with additional numerical results, are provided in the Appendix.
Throughout this paper, we adopt the following notational conventions unless otherwise specified. Scalars are denoted by regular (non-bold) letters, vectors by bold lowercase letters, and matrices by bold uppercase letters. For any positive integer $m$, let $m]=\{1,\ldots,m\}$, and for $j\in[m]$, let j-j denote the index set $m]\setminus{j}$. For a matrix $\boldsymbol{\mathbf{M}}$ and index sets $\mathcal{I}$ and $\mathcal{J}$, we write $\boldsymbol{\mathbf{M}}_{\mathcal{I},}$, $\boldsymbol{\mathbf{M}}_{\mathcal{J}}$, and $\boldsymbol{\mathbf{M}}_{\mathcal{I},\mathcal{J}}$ for the submatrices of $\boldsymbol{\mathbf{M}}$ consisting of rows indexed by $\mathcal{I}$, columns indexed by $\mathcal{J}$, and both rows indexed by $\mathcal{I}$ and columns indexed by $\mathcal{J}$, respectively.
2 Model formulation and natural parametrization
In this section, we first review the model formulation of zhangli for estimating covariate-adjusted GGMs in (1) via a nodewise regression approach, and discuss the challenges of their two-stage estimation procedure arising from the nonconvexity of the formulation. We then introduce a convex formulation of (1) based on a natural parametrization.
2.1 Nodewise regression in covariate-adjusted GGMs
Suppressing the dependence of $\boldsymbol{\mathbf{\Sigma}}$ and $\boldsymbol{\mathbf{\mu}}$ on the covariates $\bm{U}$ for notational simplicity, the Gaussian model in (1) implies that, for each $j\in[p]$,
$X_{j}-\mu_{j}\>|\>\bm{X}_{-j},\bm{U}\sim N\left(\boldsymbol{\mathbf{\Sigma}}_{j,-j}\boldsymbol{\mathbf{\Sigma}}^{-1}_{-j,-j}(\bm{X}_{-j}-\boldsymbol{\mathbf{\mu}}_{-j}),\,\Sigma_{jj}-\boldsymbol{\mathbf{\Sigma}}_{j,-j}\boldsymbol{\mathbf{\Sigma}}^{-1}_{-j,-j}\boldsymbol{\mathbf{\Sigma}}_{-j,j}\right).$ (4)
By the block matrix inversion formula, (4) is equivalent to the linear model with Gaussian errors
$X_{j}=\mu_{j}-\Omega^{-1}_{jj}\boldsymbol{\mathbf{\Omega}}_{j,-j}(\bm{X}_{-j}-\boldsymbol{\mathbf{\mu}}_{-j})+\varepsilon_{j},\quad\varepsilon_{j}\sim N(0,\sigma^{2}_{\varepsilon_{j}}),\quad\sigma^{2}_{\varepsilon_{j}}=1/\Omega_{jj},$ (5)
where the expression is understood conditional on $(\bm{X}_{-j},\bm{U})$. Based on this representation, nodewise regression methods for estimating GGMs [meinshausen] can be naturally extended to incorporate covariate adjustments, where the regression coefficients and error variance are used to recover the $j$-th column of the precision matrix $\boldsymbol{\mathbf{\Omega}}$.
2.2 Non-convexity of the nodewise regression under (3)
Write $\boldsymbol{\mathbf{\Gamma}}=(\boldsymbol{\mathbf{\gamma}}_{1},\dotsc,\boldsymbol{\mathbf{\gamma}}_{p})^{\top}\in\mathbb{R}^{p\times q}$. Under the specification (3) of zhangli, the nodewise regression model (5) becomes
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\sum_{k\neq j}^{p}\eta_{jk0}(X_{k}-\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{k})+\sum_{k\neq j}^{p}\sum_{h=1}^{q}\eta_{jkh}u_{h}(X_{k}-\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{k})+\varepsilon_{j},$ (6)
[…]
To address this difficulty, zhangli separate the estimation of $\boldsymbol{\mathbf{\gamma}}_{j}$ from that of $\eta_{jkh}$. Specifically, if $\boldsymbol{\mathbf{\Gamma}}$ were known, the responses may first be centered via $Z_{j}=X_{j}-\boldsymbol{\mathbf{u}}^{T}\boldsymbol{\mathbf{\gamma}}_{j}$, reducing the model to the oracle form
$Z_{j}=\sum_{k\neq j}^{p}\eta_{jk0}Z_{k}+\sum_{k\neq j}^{p}\sum_{h=1}^{q}\eta_{jkh}u_{h}Z_{k}+\varepsilon_{j},$ (7)
for which the least squares criterion is convex in $\eta$’s.
In practice, however, $\boldsymbol{\mathbf{\Gamma}}$ is not known, and a two-stage procedure is employed. In the first stage, each $\boldsymbol{\mathbf{\gamma}}_{j}$ is estimated under a simplified working model
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\varepsilon_{j}.$
Comparing with the full model (6), this working model imposes a strong assumption on the specification (3), effectively treating the precision matrix $\boldsymbol{\mathbf{B}}_{0}$ as diagonal and all other matrices $\boldsymbol{\mathbf{B}}_{h}$ as zero matrices, thereby assuming conditional independence among all pairs of responses, which is furthermore not dependent on covariates. As a result, the first-stage estimator does not account for the network structure and its potential dependence on covariates, and thus may incur non-negligible bias when such dependencies are present.
Given the resulting estimate of $\boldsymbol{\mathbf{\Gamma}}$, the second stage proceeds by forming centered responses $Z_{j}$ and estimating the coefficients $\eta_{jkh}$ using the oracle model (7). This stage treats the estimate of $\boldsymbol{\mathbf{\Gamma}}$ as fixed, thereby inheriting any estimation error from the first stage.
The two-stage procedure introduces additional estimation error, as inaccuracies in estimating $\boldsymbol{\mathbf{\gamma}}_{j}$ propagate to the estimation of $\eta_{jkh}$. As shown in the theoretical analysis of zhangli, the concentration guarantees are weakened relative to the oracle setting in which $\boldsymbol{\mathbf{\Gamma}}$ is known, since the overall error probability depends on the tuning parameters from the first stage. We revisit these limitations in Section 4.1.
2.3 A convex formulation via natural parametrization
[…]